OpenStudy (anonymous):

Determine 2 linearly independent power series solutions to the following differential equation: y'' - 2xy' - 2y = 0 I'm having trouble rewriting the diff. eqn. as a single sum of index 0.

OpenStudy (anonymous):

what class is it for?

OpenStudy (anonymous):

Linear Algebra/Diff Eqns

OpenStudy (anonymous):

oh , I only know the latter

OpenStudy (anonymous):

Oh, there's no linear algebra involved in solving this. I just need help rewriting the diff equation in terms of a single sum. Whenever I try to combine the 3 sums you end up with, I keep getting and excessive (x) term. Messes up the whole thing

OpenStudy (anonymous):

so do you solve it normal way first then write the soloution as power series?

OpenStudy (anonymous):

Here's how they do it in the book. I've highlighted the part that I'm having trouble with.

OpenStudy (anonymous):

i have no idea what this is about, but i can follow the algebra ok. what is the question?

OpenStudy (anonymous):

Here's how they do it in the book. I've highlighted the part that I'm having trouble with.

OpenStudy (anonymous):

Combining the 3 sums they come up with is not making sense to me. In particular the 2nd sum, when re-indexed about 0 should have an x^(k+1) term in it, but that disappears in the simplificaiton

OpenStudy (anonymous):

let me take a closer look because it seems ok to me.

OpenStudy (anonymous):

oh ok i see what the problem is, but i don't think they are reindexing (if that is a word) the second sum. it is already in terms of k and it is \[-2\sum_{k=1}^{\infty}ka_kx^k\] so at \[k=0\] this term vanishes

OpenStudy (anonymous):

clear right? because of the \[ka_k\]

OpenStudy (anonymous):

so you might as well start at 0, because at 0 this term is gone. then the algebra should make sense

OpenStudy (anonymous):

forgetting about the unimportant -2 for a second what this is saying is that \[\sum_{k=1}^{\infty}ka_kx^k+\sum_{k=0}^{\infty}a_kx^k \] \[=\sum_{k=0}^{\infty}(k+1)a_kx^k\] and you can check that it is right by replacing k by 0

OpenStudy (anonymous):

i can see why this is confusing because they did reindex the first term making \[k=n-2\]

OpenStudy (anonymous):

Here's how they do it in the book. I've highlighted the part that I'm having trouble with.

OpenStudy (anonymous):

Oh! Ha! That works! My only hiccup is that shouldn't I still be able to re-index the sum and end up with the same answer?

OpenStudy (anonymous):

you are just trying to make the exponents match up right? they match in the second and third summation, just not the first

OpenStudy (anonymous):

Here's how they do it in the book. I've highlighted the part that I'm having trouble with.

OpenStudy (anonymous):

Right, and by adding zero to the second sum, we can start at k = 0 instead of k = 1. I guess I'm just curious as to why we can't substitute k = x + 1 and end up with the exact same answer. (ie: the method we used in the first sum to end up with the proper exponents)

OpenStudy (anonymous):

i guess the best way to see that is to make the replacement and see if you get the exact same thing. i assume you mean substitute \[k=j+1\] or something right?

OpenStudy (anonymous):

Here's how they do it in the book. I've highlighted the part that I'm having trouble with.

OpenStudy (anonymous):

Right, I mean right off the bat you can see that you'll end up with an x^(k+1) if you do that. I guess it's just a matter of recognizing that starting that particular sum at zero is exactly the same as starting it at 1.

OpenStudy (anonymous):

the first one starts at n = 2 and the exponent is n - 2 so the natural thing to do is shift up by two so the exponent starts at 0 and the index starts at 0 as well